OpenStudy (anonymous):

A bag contains 10 blue, 6 red, and 4 green rubber balls. If you select a ball at random from the bag, what is the probability of not choosing a green ball? Explain why it is reasonable to eliminate choice D, even though the denominator of choice D is appropriate. I know that the answer would be 4/5 but why is it reasonable to eliminate D?

OpenStudy (anonymous):

And what is choice D?

OpenStudy (anonymous):

9/10

OpenStudy (anonymous):

Sorry, forgot to add that.

OpenStudy (zarkon):

how many balls are not green?

OpenStudy (anonymous):

16

OpenStudy (zarkon):

how many balls are there in total?

OpenStudy (anonymous):

20

OpenStudy (zarkon):

now divide the two numbers

OpenStudy (anonymous):

Divide 16 by 20?

OpenStudy (anonymous):

Or each number by 4?

OpenStudy (zarkon):

16 divided by 20

OpenStudy (anonymous):

Which equals 0.8.

OpenStudy (zarkon):

\[\frac{16}{20}=\frac{8}{10}=\frac{4}{5}=.8\]

OpenStudy (anonymous):

Ok, I get it now. Thank you for your help. :)

OpenStudy (zarkon):

no problem